Night report 2021-04-03

Observers:  Priscila Pessi, Ayan Mitra
Support: Phil Wiseman, Chris Frohmaier, Tomás Müller, Marcus Toy 
TAT: Takashi Nagao, Lluís Galbany, Shubham Srivastav

SOFI Calibrations: 
Flats, specflats, specarcs 

Weather:
Start: humidity 28% , wind  4.1 m/s, seeing 0.8 
Middle: humidity 23 %, wind  4.3 m/s, seeing 1.1
End: humidity %, wind   m/s, seeing 

Comments:
We moved the rotator at the beginning of the night because we saw one of the panels was noisy.
We did image quality for EFOSC.
The telescope was closed at 4:43 UT due to clouds, no further observations were done.

 UTTarget 
 TypeSetupExposure (s)Airmass
Seeing 
Notes
23:35SN2020oczFollow UpJHK Imaging51001.10.8
 01:00 SN2021ghc Follow Up JH Imaging 3160 1.4 0.7 
01:46SN2021bnwFollow UpJHK Imaging51001.20.7 
        Switched to EFOSC2

Switch to EFOSC2

Comments:

 UT Target Type Setup Slit Exposure[s] Airmass Seeing["] Notes
03:16AT2021gqmClassification  Gr131.010001.20.6type Ia,  z=0.086
 03:36     FOCUS       
03:44SN2021fmuRe-ClassificationGr13   120001.10.8We had an earthquake [richter 4.7] at the end of the integration

Comments